LgswUnitDx operation

8:00: we start opertations at horizon to fix the CalUnitSystem DiodeSource problem. It seems related to the low temperature in the C-ring rack: JZ puts some tape on the unit to reduce the cooling ventilation. Now in operating condition the temperature is about 20°C.

9:45: Gilles goes up to swap fibers (the fiber splitter is not installed yet)

TN: 20150221_165046 bad CalUnitSystem spots on PatCams

Then we center the spots on the PatCams and we reduce coma:

HEXA Pos
X -1.52
Y -5.57
Z 4.25
Rx -262
Ry 155
10:30: we find a misconfiguration problem in the Lgsw network, we could not talk with ASM. Fixed: DX ASM IP is 192.168.11.12 (adsecdx)

We load subapertures for TCoG algo. REC LGSW only to center the hexapod and get the reference positions: 20141202_125700

Diodes currents are set to 250, 150, 200mA to get a 4000ADU/subap fluxes.

12:00: ??? maybe PBL with arbitrator ???

We recenter again manually:

HEXA Pos
X -2.52
Y -3.77
Z 4.3
Rx -470
Ry 55
We try to close just on LGSW. We have the new Hexa coordinates:

HEXA Pos
X -2.52
Y -3.77
Z 4.35
Rx -471
Ry 56

TN Notes Plot
20150221_190749 NO disturb, LGSw only, G=0.1,0.1,0.1 20150221_190749-20150221_192500_modalplot.png
20150221_190913 G=0.5,0.3,0.2
20150221_191120 no loop, to check Snapshot entries
20150221_191549 disturb=1.2", G=0.6,0.5,0.4
20150221_191718 disturb=1.2", G=0.5,0.3,0.2
20150221_192500 disturb=1.2", G=0.7,0.7,0.6
We pause the loop and we center the FLAO board on the on-axis source.

FLAO stage Pos
X -73.78
Y -53.78
Z 42.00
We load the Combined REC (FLAO BIN2 + LGSW) TCoG and we close the loop:

We are having PBL with the CalWhiteLightSource: it disconnects and resets to closed shutter + power off positions...

TN Notes Plot
20150221_200425 Disturb=1.2", G=2.0,0.5,0.4, no TS  
20150221_210945 20150221_210945-20150221_211135_modalplot.png
20150221_211019
20150221_211035
20150221_211122 OL
20150221_211135
We spend the afternoon in trying to debug the Truth Sensing software.

Flow meter installation

We installed a flowmeter on the LgswUnitDx cable cover. It allows to tune the flow from the filter installed on the C-ring to the LGSW itself. We regulated it to get a 2l/min flux on the LGSW.

IMG_20150221_151954261.jpg

Calibration light sources

The DX green off-axis LED source was changed in the morning to a spare one (#6 -> #3). It could not be remotely operated anymore. However, the new one showed similar behavior. It was concluded that it is probably temperature. The lower part of the unit was covered with tape to avoid the cold air flowing directly through from the van below.

As mentioned above the white lite source showed some disconnects, which triggered to close the shutter after reconnection and also some flickering of the source, which looks like unstable power supply. Has to be monitored further!

Laser system test

Laser system was switched on and power was measured short after.

SX 16.82 17.31 15.47

DX 17.6 16.6 15.46

With the Arbitrator test in the late afternoon/early evening the laser have been switch on and of again.

DX laser 3 (Head 5) didn't open the internal shutter. After manual closing and opening with Ranger the laser was working properly.

LUCI support

LUCI motors have not been available through software. JZ supported the efforts to resolve the problem. A power supply was broken. After replacing with the spare there have still been some issue and not everything working. First switch cables but maybe still more. Open Issue.

-- MarcoBonaglia - 21 Feb 2015
Topic revision: r4 - 22 Feb 2015, WolfgangGaessler
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback